Details.

Monitoring the actuator for large forces; Actuators #468, #467.

11:50. Setting up for Luci.

12:02 osco target

luci stuck after loop closed.

skip first.

luci.0004,0005

SR 68%

12:33 AM

Bin3 offset test.

gain:

_073847

TN:#_074330,_

ncpa0N

44:48.502597| MAIN > arbFunc doResumeLoop done.
WfsArbitrator_94866|DEB| 693|2022-02-14 07:44:50.381580| MAIN > arbFunc updateNcpaInternal (, 'FLAOWFS', 'LUCIFER', 305.69799999999998, 0)
WfsArbitrator_94866|DEB| 694|2022-02-14 07:44:50.408704| MAIN > arbFunc updateNcpaInternal done.

TN:_074504,_074543

_074652,_074837

_small elongation,.. _074935 hope to catch this TN,...

_ MAIN > offsetXY(): starting offset of -1.544, -0.214 mm

_075930,_080041

Stop EL33

The movie of the jumps in the images on Luci , it seems that the nd30 mask moved (jumps). Barry updated IT.

1:03am Bin3.

luci is waiting and the AO prepared.

Some issues with the GCS.no image in the SH.

Abort preset and send again.

WFS setup: bin3 300 Hz.

double star

mag 16.3

_082039,_082248,_082426,_082616

luci frames: 35,36.

1:27 AM stop

No guide star in the CCD47.

Stop this one.

MOON TEST

preset to Bin1.

11 degrees from the moon.

_085448

pupils can not be fitted int eh circles, and the center pupil failed for the acquisitions.

New preset.

Luci stuck in waiting for collimation. Stop IIF preset. Nothing to stop.

Strange that it can not be cancel the preset.....

2:04 am Josh tries to authorize for Luci again....

New preset.

bin1

14.5 degrees from the moon.

091943,_092026,_092250,

luci.0039,40

A preset was sent by error from Luci and the AO executed the presetAO and clear the current preset.

A question; why did it happen before when the center pupils failed and it was necessary to clear the preset from the FLAO gui...?

seeing 1.4"

_093101,_093122,

gopt:1.25,0.2,0.13

luci.0041.

luci.0046

_093816,_093835,_093859

NCPA ON

|2022-02-14 09:40:51.496991| MAIN > arbFunc doResumeLoop done.
WfsArbitrator_94866|DEB| 2194|2022-02-14 09:41:41.872810| MAIN > arbFunc updateNcpaInternal (, 'FLAOWFS', 'LUCIFER', 0.27719100000000002, 0)
WfsArbitrator_94866|DEB| 2195|2022-02-14 09:41:42.014345| MAIN > arbFunc updateNcpaInternal done.

2-14 09:42:04.922570| MAIN > RETRY: Offset outside FoV limits. Target xyz is -137.30 8.13 52.90 (-5002) VALUE_OUT_OF_ROFFSET LARGE 100 SEC

BACK

094328,

_PAUSE ERROR TACKING TN IN PAUSE.: _094409

094525,_094727,

371|2022-02-14 09:47:57.091477| MAIN > offsetXY(): starting offset of -59.622, 73.937 mm

OUTSIDE PATROL FIELD.

_094914,

JUMP ,58.

_095132,_095227,_095255

2:55 AM STOP

FROSTY LEO.

ON-AXIS...

bin1 setup

no visible for CCD47, better to try a different WFS setup.

3:02 am ESM

configure for bino-luci

not visible in the Py.

cancel preset.

pupil rerotator out of range.

tried two angles sith the same error in positions. It can not go to those angle positions.

No light seeing in the PyWFs.

Change to 200 Hz and EMgain 600 and it doesn't see light in the Py.

preset can not be sent.

Luci waits for collimation.

IIF authorize.

idl_21074 |INF| 3847|2022-02-14 11:16:07.418031| MAIN > fsm_stop_ao end
AdSecArbitrator_21193|INF| 2933|2022-02-14 11:16:07.434334| COMMANDHANDLER > Command Stop Ao (code 2015) successfully completed

TT gains:

0.8,0.25,0.05 rms:0.63

TN:_113349,_114422

4:51 am seeing 0.77"

_115122,_115442

5:02AM

Error reseting the BCU47

|2022-02-14 12:00:37.296407| MAIN > Turning on PROG_BCU_47
WfsArbitrator_94866|INF| 3607|2022-02-14 12:00:37.296851| MAIN > Turning off RESET_BCU_47
WfsArbitrator_94866|ERR| 3608|2022-02-14 12:01:07.297870| MAIN > Cannot turn off Bcu47 RESET-Active Low. Reason: Timeout waiting for powerboard.L.RESET_BCU_47.CUR - TIMEOUT_ERROR (-1000

reset it.

This was a time out, the same one that during the power on.

5:19 am 0.84

122706,_

1.0,0.35,0.05

rms =0.7

luci script modified and it was necessary to cancel the preset.

124851,_125008,_125424,

1.05,0.35,0.05

rms0.65

6:06am seeing

_131344

6:13 seeing 0.9"

_132231,_132700

End of the night. Not more luci images.

test to get a saturation of TT, Jump the TT values to 3.85 and 5.10

not observed any saturations or skip frames.

TEST SATURATION OF TT

-132929

3.85,0.65,0.05

_133042

5.10,0.65,0.05

STOP 6:41AM

-- JuanCarlosGuerra - 14 Feb 2022
Topic revision: r3 - 14 Feb 2022, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback